# thebibliography

### v3

@@ Abstract
Effective training of Quantum Neural Networks (QNNs) is vital in the Noisy Intermediate Scale Quantum (NISQ) era, where inherent noise challenges scalability. This paper explores how quantum noise accelerates the onset of barren plateaus (BPs) and examines strategies for mitigating this effect through careful selection of qubit measurement observables. We show that BPs emerge earlier in noisy environments than in ideal conditions. However, by strategically choosing observables that aligns better with the target objective—QNNs can be trained more effectively. To this end, we study different qubit measurement observables including PauliX, PauliY, PauliZ and a customized Hermitian observable-which aligns better with the learning objective used in this paper. Our analysis covers both global and local cost functions, where global involves all qubits, and local focuses on single-qubit measurements. Our findings reveal that with a global cost function, PauliX and PauliY observables lead to flatter optimization landscapes, indicating BPs as qubit count increases, especially under noise. PauliZ sustains trainability up to $8$ qubits but encounters BPs at $10$ qubits. Notably, the customized Hermitian observable excels in noisy conditions, supporting training up to $10$ qubits. With a local cost function, PauliZ outperforms PauliX and PauliY, maintaining training efficiency up to $10$ qubits. This study underscores the importance of noise consideration in QNN training and offers a strategic approach to observable selection, enhancing QNN performance in noisy quantum environments and advancing quantum machine learning research.
@@ I Introduction
The Noisy Intermediate-Scale Quantum (NISQ) devices are an important milestone in the evolution of quantum computing. They are characterized by their intermediate scale, typically involving 50 to a few hundred qubits [1]. The NISQ devices holds the potential to solve complex problems that are intractable for classical computers, especially in optimization and simulation[2, 3]. This is due to the quantum phenomena like superposition and entanglement which they exploit. However, NISQ devices are limited by their inherent noise and error rates, which become more pronounced as the number of qubits increases[1, 4]. This noise can lead to significant challenges in maintaining quantum coherence and error correction. Despite these limitations, NISQ devices find applications in various fields including cryptography [5], drug discovery [6], financial modeling [7] and quantum machine learning [8, 9].
Variational Quantum Algorithms (VQAs) are a class of quantum algorithms that have emerged as a promising approach for leveraging the capabilities of NISQ devices [10]. They are hybrid algorithms, combining quantum and classical computation, primarily used for solving optimization and simulation problems. The core idea of VQAs is to use a quantum computer to prepare quantum states that encode potential solutions to a problem, and a classical computer to optimize the parameters of these quantum states.
Quantum Neural Networks (QNNs) are a specific application of VQAs aimed at mimicking the behavior of classical neural networks using quantum systems [10, 11]. They utilize variational Quantum Circuits (VQCs) with parameters that are adjusted during training to perform tasks such as classification, regression, or pattern recognition [8, 12, 13]. Despite the growing interest in QNNs, their scalability, a key factor for their practical utility, is significantly hindered by challenges such as the presence Barren Plateaus (BPs) in their optimization landscapes a.k.a cost function landscapes  [^1] We interchangebaly use cost function landscape and optimization landscape throughout the paper [14]. BPs are characterized by an exponential decay in the variance of parameter gradients as the number of qubits increases, which severely hinders effective optimization. This issue presents a critical obstacle in the advancement of QNNs, necessitating urgent and focused research efforts.
@@
Various strategies have been proposed to mitigate the effect of BPs [15, 16, 17]. Additionally, the underlying causes of BPs have also been extensively analyzed. Factors such as the type and amount of entanglement [18], expressibility of the QNN[19], and the globality of the cost function [20, 21] and hardware noise [22] have been identified as significant contributors to the occurrence of BPs. Among these, quantum noise stands out as a particularly pivotal factor in the NISQ era. Unlike other sources of BPs, noise is an inherent and inescapable aspect of current quantum technologies, making it critical to address, for enhancing the trainability of QNNs. As the quantum computing field progresses, addressing noise-related challenges in QNNs will be essential for realizing their full potential in practical applications.
@@ I-A Motivational Example
The robustness of QNNs against the noise in NISQ devices is a topic of considerable interest. To further elaborate the fact that noise can induce the BPs in QNNs, potentially constraining their practical utility, in Fig. 1 we show the impact of noise on the cost function landscape, which subsequently influences the training efficacy of QNNs. For the motivational analysis, a 4-qubit QNN is employed, with each qubit undergoing $R_{x}(\theta)$ and $R_{y}(\theta)$ rotation gates. Additionally, controlled-Z (CZ) gate is used to entangle the adjacent qubits. The quantum layers have an overall depth of 10, accounting to a total of 80 single-qubit and 40 two-qubit gates. The qubit measurement is performed in PauliZ measurement basis. The QNN is then trained to learn an identity function, utilizing the Adam optimizer with a learning rate of 0.1 over 50 iterations.
Under ideal, noise-free conditions, the cost function landscape exhibits significant dynamism (Fig. 1a), with multiple regions converging towards a solution. Such a landscape facilitates the optimizer’s navigation towards an optimal solution. Conversely, in the presence of noise, affecting each qubit in every layer, the cost function landscape becomes significantly flatter (Fig. 1b). This flattening renders the landscape less conducive for the optimizer, hampering its ability to locate the solution efficiently. These differences are prominently reflected in training results, where training in the absence of noise yields significantly superior outcomes compared to training under noisy conditions, as shown in Fig. 1c. Therefore, addressing the noise is crucial for the practical realization of scalable QNNs. This paper focuses on a comprehensive analysis of noise, exploring conditions under which noise may be beneficial and strategies to mitigate its impact to a feasible extent.
@@ I-B Our Contributions
We conduct a comprehensive investigation into how quantum noise affects the training dynamics of QNNs with different qubit observables. We studied the occurrence of BPs by analyzing training dynamics under noisy and noise-free conditions while keeping quantum layers’ depth constant and increasing the number of qubits. We examined both global (all qubits measured) and local (single qubit measured) cost function definitions.
  • Our analysis shows that quantum noise precipitates the emergence of BPs in QNNs with a global cost function definition, a phenomenon which typically emerges at a relatively higher qubit count in noise-free environments. However, we find that choosing the right qubit measurement observable aligning with the cost function requirements and desired quantum circuit outputs, can significantly mitigate or exploit this effect. To this end we explore various qubit measurement strategies namely; the standard PauliZ, PauliX, PauliY, and a specially designed Hermitian observable.
  • In the global cost function setting, we observe that the optimization landscapes for PauliX and PauliY become significantly flatter as the number of qubits increases, a sign of BP emergence, and is more pronounced in noisy environments. PauliZ, however, maintains satisfactory trainability up to 6 qubits and moderate at 8 qubits in noisy conditions, but encounters BPs at 10 qubits.
  • A notable finding is that QNNs employing the customized Hermitian observable with a global cost function benefit from noise, resulting in a truncated optimization landscape that eases progression towards the solution. This observable proves effective for training up to 10 qubits, demonstrating that appropriate observable selection can leverage noise to enhance QNN trainability.
  • Contrary to the global cost function, the PauliZ observable with a local cost function, shows considerable advantage under noise, maintaining efficient training up to 10 qubits showing the potential of effective training in noisy conditions. Interestingly, there is a minimal influence of the local cost function on the optimization landscapes of PauliX and PauliY observables. The predominantly flat landscapes observed for these observables under both global and local cost functions severely limit their training effectiveness.
Our findings demonstrate that appropriate observable selection can not only mitigate the detrimental effects of noise but also leverage them to enhance the trainability of QNNs.
@@ II-A Qubit Measurement and Observables
The PauliZ observable, often represented by the symbol $\sigma_{z}$ or $Z$ is one of the three Pauli matrices used in quantum mechanics. It is particularly significant for its role in describing the state of a qubit, the fundamental unit of quantum information. The PauliZ observable is a $2\times 2$ matrix represented as:
$\sigma_{z}=Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
This matrix operates on two-dimensional complex Hilbert spaces, which are the state spaces of qubits. In the context of quantum mechanics, the PauliZ observable corresponds to the measurement of spin or polarization along the Z-axis. When applied to a qubit, it differentiates between the two basis states, typically denoted as |$\ket{0}$ and |$\ket{1}$. The eigenvalues of this matrix are $+1$ and $-1$, corresponding to these basis states. A measurement resulting in +1 indicates the system is in the |$\ket{0}$ state, and $-1$ indicates the |$\ket{1}$ state.
@@
The PauliX observable, denoted by the symbol $\sigma_{x}$ or $X$ is the second Pauli matrix and is represented as:
$\sigma_{x}=X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$
The Pauli-X observable corresponds to the measurement of spin or polarization along the X-axis. When applied to a quantum state, it effectively flips the state between the two basis states, |$\ket{0}$ and |$\ket{1}$. The eigenvalues of this matrix are $+1$ and $-1$, corresponding to these flipped states. The PauliX gate is the quantum equivalent of the classical NOT gate. It flips the state of a qubit from |$\ket{0}$ to |$\ket{1}$ and vice versa, making it fundamental for qubit state manipulation. The eigenstates of the PauliX observable are not the computational basis states (|$\ket{0}$ and |$\ket{1}$), but rather the superposition states |$\frac{\ket{0}+\ket{1}}{\sqrt{2}}$ and |$\frac{\ket{0}-\ket{1}}{\sqrt{2}}$. These states are stable under the action of the PauliX observable.
@@
Hermitian observable in quantum mechanics is a fundamental concept that extends beyond the specific examples of the Pauli matrices as discussed above. A Hermitian observable is represented by a Hermitian operator, which is a linear operator $\hat{A}$ on a Hilbert space that equals its own Hermitian conjugate (or adjoint), denoted as $\hat{A}^{\dagger}$. This is mathematically expressed as:
$\hat{A}=\hat{A}^{\dagger}$
This means that for any two vectors |$\ket{a}$ and |$\ket{b}$ in the Hilbert space, the inner product $\bra{a}\hat{A}\ket{b}$ equals the complex conjugate of $\bra{b}\hat{A}\ket{a}$. In quantum mechanics, observables correspond to physical quantities that can be measured, such as position, momentum, spin, etc. The Hermitian nature of these observables ensures that the eigenvalues, which represent the possible outcomes of measurements, are real numbers. This is essential because physical measurements yield real values. The eigenvalues of a Hermitian operator are real, and the eigenvectors corresponding to different eigenvalues are orthogonal. This property is crucial in quantum mechanics, as it allows for the definition of quantum states and the prediction of measurement outcomes.
Figure 2: Detailed methodology highlighting key steps for the analysis of noise impact on the trainability of QNNs with different qubit measurement strategies. The quantum circuits used in QNN design are constructed with 4 to 10 qubits. Each qubit has RX and RY gates applied on it, and entanglement was achieved between neighboring qubits via the CZ gate. Two distinct quantum circuit configurations were explored, one where all qubits were measured, and another with only a single qubit measurement. The evaluation metrics used are optimization landscapes and cost function convergence.
@@ II-B Amplitude Damping
Amplitude damping is a concept primarily used in quantum mechanics and refers to a type of quantum noise that causes a quantum system to lose energy. Amplitude damping is a significant concern because it can lead to errors in quantum calculations. It describes how a quantum system, such as a qubit, interacts with its environment, leading to a loss of energy. This interaction can cause changes in the state populations of the qubit, and it is the underlying mechanism behind phenomena such as scattering, dissipation, attenuation, and spontaneous emission.
The mathematical description of the amplitude damping channel is typically provided by Kraus matrices. These matrices are used in quantum mechanics to represent the evolution of quantum states in open quantum systems, i.e., systems that interact with their environment. The amplitude damping channel has two primary Kraus matrices, $A_{0}$ and $A_{1}$, which are defined as follows:
@@
    $A_{0}$: This matrix describes the probability of the qubit remaining in its current state and is represented as:
    $A_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}$
    where $\gamma$ represents the damping parameter, a real number between 0 and 1, which quantifies the probability of the qubit transitioning from the higher energy state to the lower energy state.
@@
    $A_{1}$: This matrix represents the transition of the qubit from its higher energy state to its lower energy state due to the interaction with the environment and is described as:
    $A_{1}=\begin{pmatrix}0&\sqrt{\gamma}\\ 0&0\end{pmatrix}$
    This matrix captures the essence of amplitude damping: the qubit loses energy, moving from the excited (|$\ket{1}$) to ground (|$\ket{0}$) state with a probability proportional to $\gamma$.
@@ III HQNET Methodology
In this paper, we investigate the influence of noise on the training dynamics of QNNs. Our analysis primarily utilizes the hardware-efficient ansatz, a popular approach in NISQ applications characterized by sequences of single and two-qubit gates. The general form of these circuits is given by the following Equation:
@@
where $U_{\text{ent}}$ represents a two-qubit gate(s) typically used for qubit entanglement and $U_{\text{\text{rotation}}}(\theta_{i})$ represents a single qubit parameterized gate, whose parameters are optimized during training, and $N$ is the total number of repetitions of the quantum circuit until measurement.
We focus on a comprehensive analysis of the cost function or optimization landscape of QNNs under both noisy and ideal, noise-free scenarios, considering a range of qubit measurement strategies. A key component of our approach involves incrementally increasing the number of qubits, which is crucial for understanding the influence of noise in the context of Barren Plateaus (BP). The concept of BP suggests that the cost function landscape in QNNs tends to flatten with an increase in the number of qubits, thereby posing a significant challenge for optimizers in locating the optimal solution. Moreover, our exploration delves into how different qubit measurement strategies, when applied in noisy conditions, affect the trainability of QNNs. A pivotal aspect of this investigation is to ascertain whether noise can be strategically utilized to enhance the learning process in QNN training. By systematically varying the parameters and conditions, we aim to uncover insights into the role of noise and its potential benefits in the optimization of QNNs. A detailed overview of our methodology is depicted in Fig. 2.

### v4

@@ Abstract
Effective training of Quantum Neural Networks (QNNs) is vital in the Noisy Intermediate Scale Quantum (NISQ) era, where inherent noise challenges scalability. This paper explores how quantum noise accelerates the onset of barren plateaus (BPs) and examines strategies for mitigating this effect through careful selection of qubit measurement observables. We demonstrate that BPs emerge earlier in noisy environments than in ideal conditions. However, by strategically choosing observables that aligns better with the target objective, QNNs can be trained more effectively. To this end, we study different qubit measurement observables including PauliX, PauliY, PauliZ and a customized Hermitian observable-which aligns better with the learning objective used in this paper. Our analysis covers both global and local cost functions, where global involves all qubits, and local focuses on single-qubit measurements. Our findings reveal that with a global cost function, PauliX and PauliY observables lead to flatter optimization landscapes, indicating BPs as qubit count increases, especially under noise. PauliZ sustains trainability up to $8$ qubits but encounters BPs at $10$ qubits. Notably, the customized Hermitian observable excels in noisy conditions, supporting training up to $10$ qubits. With a local cost function, PauliZ outperforms PauliX and PauliY, maintaining training efficiency up to $10$ qubits. This study underscores the importance of noise consideration in QNN training and offers a strategic approach to observable selection, enhancing QNN performance in noisy quantum environments and advancing quantum machine learning research.
@@ I Introduction
The advent of Noisy Intermediate-Scale Quantum (NISQ) devices—quantum processors with 50 to a few hundred qubits—marks a significant step toward practical quantum computing [1]. By harnessing superposition and entanglement, NISQ devices offer potential advantages in solving classically intractable problems in optimization and simulation [2, 3]. However, their utility is constrained by high noise levels and limited error correction, which become increasingly problematic with scale [1, 4]. Despite these challenges, NISQ devices are already showing promise in cryptography [5], drug discovery [6], financial modeling [7], and quantum machine learning [8, 9].
Variational Quantum Algorithms (VQAs) are a hybrid class of algorithms that leverage both quantum and classical computation, making them well-suited for NISQ devices [10]. They solve optimization and simulation problems by using quantum circuits to encode solutions and classical optimizers to adjust circuit parameters. Quantum Neural Networks (QNNs) are a subclass of VQAs designed to emulate classical neural networks using variational quantum circuits (VQCs) [10, 11]. These models are trained by tuning circuit parameters to perform tasks such as classification, regression, and pattern recognition [8, 12, 13]. Despite their potential, QNNs face scalability issues due to Barren Plateaus (BPs). BPs are the regions in the cost function landscape where gradient variance vanishes exponentially with increasing qubit count [14]. This severely impedes optimization and poses a major barrier to the practical deployment of QNNs, necessitating focused research on mitigating BPs.
@@
Several strategies have been proposed to mitigate BPs [15, 16, 17], alongside efforts to understand their underlying causes. Key contributing factors include the nature and degree of entanglement [18], circuit expressibility [19], cost function globality [20, 21], and hardware-induced noise [22]. Among these, quantum noise is particularly critical in the NISQ era, as it is inherent to current quantum devices and directly impacts the trainability of QNNs. Addressing noise-related challenges is thus essential for advancing the practical deployment of QNNs.
@@ I-A Motivational Example
The robustness of QNNs to noise in NISQ devices remains a critical concern. To highlight how noise can induce Barren Plateaus (BPs) and hinder trainability, Fig. 1 illustrates its impact on the cost function landscape. We employ a $4$-qubit QNN composed of alternating $R_{x}(\theta)$ and $R_{y}(\theta)$ rotations, with adjacent qubits entangled via controlled-Z (CZ) gates. The circuit depth is $10$, comprising $80$ single-qubit and 40 two-qubit gates, with measurements in the Pauli-Z basis. The model is trained to learn the identity function using the Adam optimizer (learning rate = $0.1$) for $50$ iterations. Under ideal conditions, the cost landscape (Fig. 1(a)) exhibits a rich structure and multiple minima, enabling efficient optimization. However, in noisy settings where each qubit is affected in every layer, the landscape flattens significantly (Fig. 1(b)), impeding convergence. This degradation is evident in the training performance (Fig. 1(c)), where noise-free training substantially outperforms the noisy counterpart. These findings underscore the importance of addressing noise to realize scalable QNNs. This paper presents a detailed analysis of noise-induced effects, identifying conditions where noise may be leveraged and proposing strategies for its mitigation.
@@ I-B Our Contributions
We investigate how quantum noise and observable choice affect QNN training dynamics and barren plateaus, comparing global vs. local cost functions under fixed-depth circuits with increasing qubits. Our contributions are summarized below:
  • Our analysis reveals that quantum noise accelerates the onset of BPs in QNNs with global cost function definitions—an effect that typically occurs at higher qubit counts under noise-free conditions. However, we demonstrate that selecting appropriate qubit measurement observables, aligned with the cost function and target circuit output, can mitigate or even exploit this effect. To this end, we evaluate multiple measurement strategies, including PauliZ, PauliX, PauliY, and a custom-designed Hermitian observable.
  • In the global cost function setting, we observe that the optimization landscapes for PauliX and PauliY measurements become increasingly flat as the number of qubits increases, which is a clear indication of BPs, and is further exacerbated by noise. In contrast, PauliZ measurement maintains performance up to $6$ qubits and shows moderate degradation at $8$ qubits under noisy conditions, with BPs becoming evident at $10$ qubits (maximum width considered in this paper).
  • A notable finding is that QNNs utilizing the customized Hermitian observable with a global cost function exhibit noise-induced benefits, resulting in a truncated yet structured optimization landscape that facilitates convergence. This observable enables effective training up to $10$ qubits, highlighting that careful observable selection can harness noise to improve QNN trainability.
  • In contrast to the global cost function, the PauliZ observable under a local cost function demonstrates strong resilience to noise, enabling efficient training up to $10$ qubits and underscoring its suitability for noisy environments. Interestingly, PauliX and PauliY exhibit minimal improvement under the local cost setting, with consistently flat optimization landscapes that persist across both global and local definitions, severely limiting their trainability.
Our findings demonstrate that appropriate observable selection can not only mitigate the detrimental effects of noise but can also leverage the noise to enhance the trainability of QNNs.
@@ II-A Qubit Measurement and Observables
The PauliZ observable, denoted by $\sigma_{z}$ or $Z$, is one of the three fundamental Pauli matrices in quantum mechanics. It plays a crucial role in describing the state of a qubit. The PauliZ matrix is defined as:
$\sigma_{z}=Z=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ (1)
This $2\times 2$ Hermitian matrix acts on the two-dimensional complex Hilbert space of a single qubit. In quantum mechanics, $\sigma_{z}$ corresponds to measurements along the Z-axis, distinguishing between the computational basis states |$\ket{0}$ and |$\ket{1}$. Its eigenvalues are $+1$ and $-1$, with $+1$ associated with |$\ket{0}$ and $-1$ with |$\ket{1}$. Thus, a measurement in the PauliZ basis reveals whether the qubit is aligned (|$\ket{0}$) or anti-aligned (|$\ket{1}$) with the Z-axis.
@@
The PauliX observable, denoted by $\sigma_{x}$ or $X$, is the second of the three Pauli matrices and is defined as:
$\sigma_{x}=X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$ (2)
This observable corresponds to measurements along the X-axis in the Bloch sphere representation. When applied to a qubit, it acts as a bit-flip operator, exchanging the computational basis states |$\ket{0}$ and |$\ket{1}$. The PauliX gate is thus analogous to the classical NOT gate, making it fundamental for qubit state manipulation. The eigenvalues of $\sigma_{x}$ are $+1$ and $-1$, but its eigenstates differ from the standard basis. Instead, they are the symmetric and anti-symmetric superposition states:
[…]
These eigenstates remain invariant under the action of $\sigma_{x}$, making them the stable states for measurements in the PauliX basis.
@@
A Hermitian observable in quantum mechanics refers to any observable represented by a Hermitian operator, denoted by $\hat{A}$. A Hermitian operator is a linear operator on a Hilbert space that is equal to its own Hermitian conjugate (or adjoint), i.e.,
$\hat{A}=\hat{A}^{\dagger}$ (3)
This implies that for any two vectors |$\ket{a}$ and |$\ket{b}$ in the Hilbert space, the inner product satisfies:
[…]
In quantum mechanics, Hermitian operators are associated with physical observables such as position, momentum, spin, and energy. The Hermitian property guarantees that all eigenvalues of $\hat{A}$, which correspond to the possible outcomes of a measurement are real. Moreover, eigenvectors associated with distinct eigenvalues are orthogonal, a feature that underpins the probabilistic interpretation of quantum states and the predictability of measurement results.
This general framework encompasses well-known observables like the Pauli matrices, but also allows for the design of custom Hermitian observables tailored to specific tasks in quantum algorithms or quantum neural network training.
Figure 2: Detailed methodology highlighting key steps for the analysis of noise impact on the trainability of QNNs with different qubit measurement strategies. The quantum circuits used in QNN design are constructed with 4 to 10 qubits. Each qubit has RX and RY gates applied on it, and neighboring qubits are entangled via the CZ gate. Two distinct quantum circuit configurations were explored, Global QNN (all qubits were measured), and Local QNN (only a single qubit measured). The evaluation metrics used are optimization landscapes and cost function convergence.
@@ II-B Amplitude Damping
Amplitude damping is a fundamental quantum noise model that describes energy dissipation in a quantum system, such as a qubit. It captures how interaction with the environment leads to energy loss, resulting in state transitions from excited to ground states. This process underlies physical phenomena such as spontaneous emission, scattering, and attenuation, and poses a significant challenge for reliable quantum computation.
The evolution of a quantum system under amplitude damping is commonly described using Kraus operators, which characterize the dynamics of open quantum systems. The amplitude damping channel is represented by two Kraus matrices, $A_{0}$ and $A_{1}$, defined as:
@@
    $A_{0}$ — Describes the probability of the qubit remaining in its current state:
    $A_{0}=\begin{pmatrix}1&0\\ 0&\sqrt{1-\gamma}\end{pmatrix}$ (4)
    where $\gamma\in[0,1]$ is the damping parameter representing the probability of energy loss.
@@
    $A_{1}$ — Represents the transition of the qubit from the excited state to the ground state:
    $A_{1}=\begin{pmatrix}0&\sqrt{\gamma}\\ 0&0\end{pmatrix}$ (5)
    This operator captures the decay of the qubit from |$\ket{1}$ to |$\ket{0}$ with probability $\gamma$ due to environmental interaction.
Together, these Kraus operators satisfy the completeness relation $A_{0}^{\dagger}A_{0}+A_{1}^{\dagger}A_{1}=I$, ensuring that the quantum channel preserves trace and thus remains physically valid.
@@ III HQNET Methodology
We investigate the impact of quantum noise on the training dynamics of Quantum Neural Networks (QNNs), employing the hardware-efficient ansatz commonly used in NISQ-era applications. These circuits consist of parameterized single-qubit rotations and fixed two-qubit entangling gates, represented as:
@@
where $U_{\text{ent}}$ denotes the entangling gates, $U_{\text{rot}}(\theta_{i})$ represents trainable single-qubit rotation gates, and $N$ is the number of circuit layers. We conduct a comprehensive analysis to assess whether the quantum noise which is typically regarded as detrimental, can be strategically leveraged to improve optimization. Our methodology, illustrated in Fig. 2, systematically varies the number of qubits and measurement observables to uncover insights into the role of noise in QNN training.